OpenStudy (anonymous):

Check my answers? :) And #5 please.

ganeshie8 (ganeshie8):

remember this : 'arc measure' is same as 'central angle'

ganeshie8 (ganeshie8):

#2 : \(\angle FDE = 78/2 = 39\)

ganeshie8 (ganeshie8):

#1 adn #3 are correct

OpenStudy (anonymous):

oh, right... got confused .__.

ganeshie8 (ganeshie8):

tell me what #4 shoudl be now ?

OpenStudy (anonymous):

102?

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

i got confused with #5 too

ganeshie8 (ganeshie8):

for #5 simply use one of below :- 1) triangle sum property 2) triangle exterior angle theorem

OpenStudy (anonymous):

right.. :) i didnt see that i ady found D xD

ganeshie8 (ganeshie8):

good :) wat about #6

OpenStudy (anonymous):

156 :)

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

258?

ganeshie8 (ganeshie8):

Yes :) 360 - 102 = 258

OpenStudy (anonymous):

yayyyy thanks c:

ganeshie8 (ganeshie8):

good job !!

Directrix (directrix):

@stupidinmath Will you check my work?

OpenStudy (anonymous):

sure2 :)

OpenStudy (anonymous):

36? I think there's some typo @Directrix :)) But yeah, therest, we got same :DD

Directrix (directrix):

Let me look.

Directrix (directrix):

@stupidinmath What did you get for <FDO?

OpenStudy (anonymous):

39..

Directrix (directrix):

Oh, m<FDO should be (1/2) of 78 which is 39. My mistake. Thanks for checking.

OpenStudy (anonymous):

yeah, thank you too :)